# Forecasting the future development in quality and value of professional football players

### v2

@@ 2.2.2 Player value
In contrast with player quality, there does exist a ground truth for the concept of player values. The value of a transfer fee is based on the value of the player for each of the involved clubs and historical transfers of similar players [37]. Thus, the historical values of these fees can be used to predict the transfer value of a player, albeit at the cost of some selection bias.
@@
A well-known medium that describes the value of a player is Transfermarkt. This company uses crowd estimation to assign values to players [38]. These market values describe the general value of a player and do not take into account the temporary situation of a football player, like the current club and contract length. This means that they describe a different quantity than the expected value of a transfer fee. Nonetheless, these values are strongly correlated with the real transfer fees as shown by Herm et al. [38]. Therefore, both market values by Transfermarkt and transfer values are often used interchangeably when describing the monetary value of a football player.
@@
The literature review by Franceschi et al. [24] showed that many studies have been performed to describe the monetary value of a football player based on data. The authors found that most of these studies used linear regression to find which variables have a significant linear dependence on the value of a football player. The review by Franceschi et al. [24] considered 111 trained models that were used in the scientific literature to investigate the transfer value of a football player. The vast majority (85%) of the models are based on ordinary least squares. The authors have also shown the importance of different variables in the considered models. For instance, they have found that age, the square of the age, and the number of matches played by a player are frequently studied variables. These variables are also most often found to be significant. Players frequently increase in value as they get better with experience that is gained over the years, but they also decrease in value as a player loses the potential to improve when getting older. Consequently, the influence of age is often measured with a quadratic term. Similarly, the number of games played can be expected to be an important variable because players gain experience by playing games, which makes them more valuable. Moreover, players who play a lot of games are often the better players on a team. This explains why these variables are important, as found by Franceschi et al. [24]. Their study also shows that almost no variables describing defensive behavior are considered when other researchers trained models to describe transfer values. As a consequence, the resulting models can be expected to describe the value of offensive players better than that of defensive players. In addition, the linear models in these studies are mostly trained to determine the influence of variables on the transfer value of a football player, and they are generally not trained and tested for out-of-sample prediction, which limits the application of predicting based on new, unseen data.
@@
In contrast, Al-Asadi and Tasdemır [39] have trained multiple models to predict market values of football players with features from the video game FIFA. They have found that a random forest model gives improved predictions over linear methods. Steve Arrul et al. [40] have studied the application of artificial neural networks for the same problem, also considering features from the video game FIFA. They obtain similar loss values as the random forest model of Al-Asadi and Tasdemır [39]. The research by Behravan and Razavi [41] introduces a methodology to train a support vector regression model via particle swarm optimization for the prediction of market values. Although the data set is somewhat similar to the studies above, this model attains worse loss values. In the study of Yang et al. [1], random forests, GAMs, and QAMs are applied to predict the transfer fees of players based on variables describing the player. The authors have inferred from their random forest model that the expenditure of the buying club and the income of the selling club are important features in predicting the transfer fee, as well as the age and the remaining contract duration. This research additionally illustrates how GAMs and QAMs can be used to investigate the dependency between the player transfer value and the given features. The QAM models show that this relation varies for different quantiles, indicating the need to study the influence of models on different groups of players. On the other hand, the GAMs show that the relationship between the transfer values and the features is often nonlinear. A study with extensive types of player performance metrics as features has been performed by McHale and Holmes [2], in which linear regression, linear mixed effects, and XGBoost models have been trained. These models not only include statistics such as the number of minutes played, height, and position, but also plus-minus ratings based on xG, expert ratings from the video game FIFA, and GIM ratings, which are similar to VAEP ratings. The results show that the best predictive performance is attained by the XGBoost model, although the linear mixed effects with the buying and selling clubs as random effects also provide good results. Their results indicate that their model outperforms Transfermarkt market values when predicting the transfer fees on average, although the market values are a better predictor for transfers of more than €20 million. These studies show how supervised learning can be used to obtain models that predict the value of a football player. They show that nonlinear methods generally predict more accurately and that the patterns can differ for different groups of players.
@@ 2.3 Predicting future values
Although the studies discussed above are concerned with the quantification of the quality and monetary value of football players at the present moment, only limited work has been done on the future development of player performance. Apostolou and Tjortjis [42] have conducted a small-scale study with the aim of predicting the number of future goals of the two football players Lionel Messi and Luis Suárez using a random forest, logistic regression, a multi-layer perception classifier, and a linear support vector classifier. Pantzalis and Tjortjis [43] have predicted the expert ratings in the next season of 59 center-backs in the English Premier League based on one season of player attributes from a popular football manager simulation game. Their method uses a linear regression model to describe the in-sample patterns. Giannakoulas et al. [44] have trained linear regression, random forest, and multi-layer perceptron models to predict the number of goals of a football player in a season before the start of the corresponding season. Their dataset entails around 800 football players. Similarly, the models by Markopoulou et al. [45] have been trained to predict the number of goals by looking at the creation of different models per competition. This study on 424 football players shows that the best results are often obtained using XGBoost models for this prediction problem and that making different models for different competitions might be beneficial.
@@
Barron et al. [46] have tried to predict the tier within the English first three leagues in which a football player would play next season as an indicator of player quality. Three artificial neural networks are trained to predict in which league a player would play with the data of 966 football players. Their models are only able to recognize the differences between players in the lowest and highest tiers (League One and the English Premier League).
@@
Little literature exists about the prediction of the future transfer values of football players. Baouan et al. [47] have applied lasso regression and a random forest model to identify important features for the development of around 22,000 football players. They have trained these supervised models for players of different positions to predict a player’s transfer value two years in the future based on performance statistics. The feature importances of the models show, for instance, that the average market value of a league is an important feature for the future values of players in that league. Although cross-validation is performed for the hyperparameter tuning, this study focuses on finding in-sample patterns.
@@
Our research treats forecasting of both player quality and transfer value. The current paper builds on the existing literature about forecast player performance by studying the long-term forecasting of a model-based player quality indicator on a larger data set. Additionally, this research avoids the bias in current literature that better describes offensive players by using a more general top-down rating. Our research contributes to the existing knowledge of forecasting the monetary value because the models are trained to perform out-of-sample prediction, which makes it possible to apply them to unseen situations. Moreover, the combination of forecasting the development of both in player quality and transfer value gives a comprehensive summary of the most important factors in transfer decisions. In this way, this paper fills a gap in the literature by forecasting the development of model-based indicators for player quality and value on a significant data set in a predictive setting.
@@ 3.1.1 Player quality: SciSkill
The first prediction problem concerns predicting the development in the subsequent year of the player quality described by an EM algorithm called the SciSkill [23], which is a generalization of the Elo rating as discussed in section 2. The dataset is restructured to contain monthly data points describing the situation of each player at that time. Each monthly data point consists of the features and the dependent variable. The dependent variable is the difference between the player quality one year ahead and the current player quality value, which is the development in player quality in one year. The features set consist of 86 features constructed with domain knowledge describing for instance the month of year, the current player performance, league strength, time since the most recent game, player characteristics, time series information of the SciSkill, the club’s transfer situation, and the difference in quality between the player and his teammates. The dependent variable and features are described in LABEL:Appendix:_Table_dependent_variable_player_quality and LABEL:Appendix:_Table_features_player_quality.
Only male players with more than 20 games and more than 2 years of data are considered. The final data set consists of 80,568 male professional football players playing in the years 2012 to 2023. As the data set consists of 3,834,539 data points, there are on average 47.6 monthly data points per player, which corresponds to roughly 4 seasons of data.
@@ 3.1.2 Player value: Estimated Transfer Value
The data set for this prediction problem of the development in player value describes the current situation of professional football players with 58 features. These features were constructed based on domain knowledge and data availability. Because there were fewer data points than for the player quality case study, the number of features was reduced by eliminating highly correlated features. These features consist of indicators of league strength, age, experience, contract situation, the current quality (SciSkill), or the monetary value (ETV) of a football player. The feature set also includes player characteristics, like playing position and age, as well as the differences in quality with teammates, the transfer history of his current club, and league strength. A description of all features and the dependent variable in this prediction problem can be found in LABEL:Appendix:_Table_dependent_variable_player_value and LABEL:Appendix:_Table_features_player_value.
@@ 3.2.1 Model assessment
Supervised models are trained on the prediction problems of player quality and player value. The root mean square error (RMSE) and mean absolute error (MAE) are determined for the test set. The RMSE is more strongly influenced by large errors, which can be expected to happen for ‘superstar’ players. Since these players are especially of interest to practitioners, the performance in the RMSE is considered the most important. The losses are determined on different parts of the test set. First, the loss values of the RMSE and MAE are calculated using the complete test sets. Second, the loss function of the RMSE is also considered for different age groups on the test sets because estimating the potential of a player is mostly interesting for young players. Third, the RMSE is studied on important subgroups of players, like players with large positive or negative development, players with good performance, or players with a high transfer value. By determining the test losses separately on the general population of players, young players, and important subgroups of players, the models are studied on their predictive performance.
Beforehand, 5% of all football players are left out of both data sets based on stratified sampling for internal studies by the data provider. Because the data of player performance indicators is often dependent on time [26, 16], time-dependent train-test splits are applied to study the predictive performance on unseen data as visualized in Figure 2. In both case studies, all data up to 2020 is considered as the training set, and from 2021 and later as the test set.
@@ 3.2.2 Model training
In this study, linear, tree-based, and kNN-based models are implemented. The linear models are ordinary least squares (OLS), also known as multiple linear regression (MLR), lasso regression, and a linear mixed effect model (LME). For the OLS model [48], feature selection is performed by applying backwards selection with a threshold of the p-values of 0.0001 for the player quality and 0.001 for the player value. These values have been determined via trial and error, considering values of $10^{k}$ for multiple values of $k$, and they are smaller than commonly used for significance testing. This is due to the predictive nature of this study and the large data sets combined with the fact that common assumptions of the linear regression model, like normality, do not hold. For lasso regression [49], feature selection is applied by selecting only the variables with nonzero coefficients. The linear mixed effect (LME) model [48] is trained to take into account the influence of a player’s nationality as a random influence. The feature selection for the LME model is done by taking the 20 variables corresponding to the largest feature importances within the lasso model. This is done for computational feasibility. For these three linear models, no interaction effects are included.
@@
The tree-based models are the decision tree [49], random forest [49], and XGBoost [30] models. For the tree-based models, feature selection is performed by first adding noise variables to the data set and training the models. All features with a larger feature importance than the noise variables are then selected. As tree-based methods can have a bias favoring non-discrete variables, both discrete and continuous variables are added. The discrete features are compared with the discrete noise variables, and the continuous features with the continuous noise variables.
@@
To investigate the predictive power of time series, three k-nearest neighbors (kNN) models have been implemented using the Hierarchical Navigable Small Worlds indexer provided by Douze et al. [50]. This was done on an altered feature set consisting of time series information with lagged versions of the most important player indicators. Because the kNN model suffers from the curse of dimensionality [27, Section 2.5], the feature set is kept relatively small. The features and labels of this altered data set are described in LABEL:Appendix:_Table_knn_features_player_quality and LABEL:Appendix:_Table_dependent_variable_player_value in the appendix. First, a normal kNN model is applied to the predictive problem. This model searches for the most similar data points with respect to the Euclidean norm. The prediction is then obtained by applying a weighted average, where closer data points are weighted more heavily as described by Dudani [31]. Possible weighting methods are the reciprocal of the absolute value of the distance, the distance with min-max scaling, or uniform weights. The method of calculating these weights is considered a hyperparameter. The second kNN model is constructed similarly, but it calculates the distances based on the Mahalanobis distance. The Mahalanobis distance projects the features on a decorrelated feature space and calculates the Euclidean distance. This makes it possible to better distinguish differences because the lagged time series features are heavily dependent. Lastly, an adapted RReliefF method [51] is implemented to calculate feature importances of the normal kNN model in a regression context. The features are then multiplied by the feature importances before calculating Euclidean distances to introduce new feature weights. The kNN model is then trained on these reweighted features.
@@ 3.2.3 Hyperparameter tuning
To select the best hyperparameters of the models in the case studies, Bayesian optimization [52] is applied to find hyperparameters minimizing the RMSE loss. Because of the time dependencies of player performance indicators, an adjusted version of cross-validation is implemented that splits per year and incrementally grows the training set as visualized for the player quality prediction problem in Figure 3. After the hyperparameter tuning, the models are trained on the complete training set with the optimized hyperparameters.

### v3

@@ 2.2.2 Player Value
In contrast with player quality, there does exist a ground truth for the concept of player values. The value of a transfer fee is based on the value of the player for each of the involved clubs and historical transfers of similar players (Poli et al., 2022). Thus, the historical values of these fees can be used to predict the transfer value of a player, albeit at the cost of some selection bias.
@@
A well-known medium that describes the value of a player is Transfermarkt. This company uses crowd estimation to assign values to players (Herm et al., 2014). These market values describe the general value of a player and do not take into account the temporary situation of a football player, like the current club and contract length. This means that they describe a different quantity than the expected value of a transfer fee. Nonetheless, these values are strongly correlated with the real transfer fees as shown by Herm et al. (2014). Therefore, both market values by Transfermarkt and transfer values are often used interchangeably when describing the monetary value of a football player.
@@
The literature review by Franceschi et al. (2024) showed that many studies have been performed to describe the monetary value of a football player based on data. The authors found that most of these studies used linear regression to find which variables have a significant linear dependence on the value of a football player. The review by Franceschi et al. (2024) considered 111 trained models that were used in the scientific literature to investigate the transfer value of a football player. The vast majority (85%) of the models are based on ordinary least squares. The authors have also shown the importance of different variables in the considered models. For instance, they have found that age, the square of the age, and the number of matches played by a player are frequently studied variables. These variables are also most often found to be significant. Players frequently increase in value as they get better with experience that is gained over the years, but they also decrease in value as a player loses the potential to improve when getting older. Consequently, the influence of age is often measured with a quadratic term. Similarly, the number of games played can be expected to be an important variable because players gain experience by playing games, which makes them more valuable. Moreover, players who play a lot of games are often the better players on a team. This explains why these variables are important, as found by Franceschi et al. (2024). Their study also shows that almost no variables describing defensive behavior were considered when other researchers trained models to describe transfer values. As a consequence, the resulting models can be expected to describe the value of offensive players better than that of defensive players. In addition, the linear models in these studies are mostly trained to determine the influence of variables on the transfer value of a football player, and they are generally not trained and tested for out-of-sample prediction, which limits the application of predicting based on new, unseen data.
@@
In contrast, Al-Asadi and Tasdemır (2022) have trained multiple models to predict market values of football players with features from the video game FIFA. They have found that a random forest model gives improved predictions over linear methods. Steve Arrul et al. (2022) have studied the application of artificial neural networks for the same problem, also considering features from the video game FIFA. They obtain similar loss values as the random forest model of Al-Asadi and Tasdemır (2022). The research by Behravan and Razavi (2021) introduces a methodology to train a support vector regression model via particle swarm optimization for the prediction of market values. Although the data set is somewhat similar to the studies above, this model attains worse loss values. In the study of Yang et al. (2024), random forests, GAMs, and QAMs are applied to predict the transfer fees of players based on variables describing the player. The authors have inferred from their random forest model that the expenditure of the buying club and the income of the selling club are important features in predicting the transfer fee, as well as the age and the remaining contract duration. This research additionally illustrates how GAMs and QAMs can be used to investigate the dependency between the player transfer value and the given features. The QAM models show that this relation varies for different quantiles, indicating the need to study the influence of models on different groups of players. On the other hand, the GAMs show that the relationship between the transfer values and the features is often nonlinear. A study with extensive types of player performance metrics as features has been performed by McHale and Holmes (2022), in which linear regression, linear mixed effects, and XGBoost models have been trained. These models not only include statistics such as the number of minutes played, height, and position, but also plus–minus ratings based on xG, expert ratings from the video game FIFA, and GIM ratings, which are similar to VAEP ratings. The results show that the best predictive performance is attained by the XGBoost model, although the linear mixed effects with the buying and selling clubs as random effects also provide good results. Their results indicate that their model outperforms Transfermarkt market values when predicting the transfer fees on average, although the market values are a better predictor for transfers of more than EUR 20 million. These studies show how supervised learning can be used to obtain models that predict the value of a football player. They show that nonlinear methods generally predict more accurately and that the patterns can differ for different groups of players.
@@ 2.3 Predicting Future Values
Although the studies discussed above are concerned with the quantification of the quality and monetary value of football players at the present moment, only limited work has been conducted on the future development of player performance. Apostolou and Tjortjis (2019) have conducted a small-scale study with the aim of predicting the number of future goals of the two football players Lionel Messi and Luis Suárez using a random forest, logistic regression, a multi-layer perceptron classifier, and a linear support vector classifier. Pantzalis and Tjortjis (2020) have predicted the expert ratings in the next season of 59 center-backs in the English Premier League based on one season of player attributes from a popular football manager simulation game. Their method uses a linear regression model to describe the in-sample patterns. Giannakoulas et al. (2023) have trained linear regression, random forest, and multi-layer perceptron models to predict the number of goals of a football player in a season before the start of the corresponding season. Their data set entails around 800 football players. Similarly, the models by Markopoulou et al. (2024) have been trained to predict the number of goals by looking at the creation of different models per competition. This study on 424 football players shows that the best results are often obtained using XGBoost models for this prediction problem and that making different models for different competitions might be beneficial.
@@
Barron et al. (2018) have tried to predict the tier within the English first three leagues in which a football player would play next season as an indicator of player quality. Three artificial neural networks are trained to predict in which league a player would play with the data of 966 football players. Their models are only able to recognize the differences between players in the lowest and highest tiers (League One and the English Premier League).
@@
Little literature exists about the prediction of the future transfer values of football players. Baouan et al. (2022) have applied lasso regression and a random forest model to identify important features for the development of around 22,000 football players. They have trained these supervised models for players of different positions to predict a player’s transfer value two years in the future based on performance statistics. The feature importances of the models show, for instance, that the average market value of a league is an important feature for the future values of players in that league. Although cross-validation is performed for the hyperparameter tuning, this study focuses on finding in-sample patterns.
@@
Our research treats forecasting of both player quality and transfer value. The current paper builds on the existing literature about forecast player performance by studying the long-term forecasting of a model-based player quality indicator on a larger data set. Additionally, this research avoids the bias in the current literature that better describes offensive players by using a more general top-down rating. Our research contributes to the existing knowledge of forecasting the monetary value because the models are trained to perform out-of-sample prediction, which makes it possible to apply them to unseen situations. Moreover, the combination of forecasting the development of both player quality and transfer value gives a comprehensive summary of the most important aspects in transfer decisions. In this way, this paper builds upon the existing literature by forecasting the development of model-based indicators for player quality and value on a significant data set in a predictive setting.
@@ 3.1 Data
Since the used SciSkill and ETV models are proprietary, exact reproduction of the results using the same models is not possible. However, similar models exist for both the SciSkill Wolf et al. (2020) and ETV models Al-Asadi and Tasdemır (2022); Steve Arrul et al. (2022); Behravan and Razavi (2021); Yang et al. (2024); McHale and Holmes (2022), as discussed in Section 2, ensuring that the overall approach remains, in principle, reproducible.
@@ 3.1.1 Player Quality: SciSkill
The first prediction problem concerns predicting the development in the subsequent year of the player quality described by an EM algorithm called the SciSkill SciSports (2020), which is a generalization of the Elo rating, as discussed in Section 2. The data set is restructured to contain monthly data points describing the situation of each player at that time. Each monthly data point consists of the features and the dependent variable. The dependent variable is the difference between the player quality one year ahead and the current player quality value, which is the development in player quality in one year. The features set consists of 86 features constructed with domain knowledge describing, for instance, the month of year, the current player performance, league strength, time since the most recent game, player characteristics, time series information of the SciSkill, the club’s transfer situation, and the difference in quality between the player and his teammates. The dependent variable and features are described in Tables S1–S3 in the Supplementary Materials.
Only male players with more than 20 games and more than 2 years of data are considered. This filtering might introduce a bias in the data against player groups on which less data is available, such as younger or injury-prone players. However, it is necessary to obtain enough data points per player. The results on players with limited data availability should thus be interpreted with care. The final data set consists of 80,568 male professional football players playing in the years 2012 to 2023. As the data set consists of 3,834,539 data points, there are on average 47.6 monthly data points per player, which corresponds to roughly 4 seasons of data.
@@ 3.1.2 Player Value: Estimated Transfer Value
The data set for this prediction problem of the development in player value describes the current situation of professional football players with 58 features. These features were constructed based on domain knowledge and data availability. Because there were fewer data points than for the player quality case study, the number of features was reduced by eliminating highly correlated features. These features consist of indicators of league strength, age, experience, contract situation, the current quality (SciSkill), or the monetary value (ETV) of a football player. The feature set also includes player characteristics, like playing position and age, as well as the differences in quality with teammates, the transfer history of his current club, and league strength. A description of all features and the dependent variable in this prediction problem can be found in Tables S4–S6 in the Supplementary Materials.
@@ 3.2.1 Model Assessment
Supervised models are trained on the prediction problems of player quality and player value. The root mean square error (RMSE) and mean absolute error (MAE) are determined for the test set. The RMSE is more strongly influenced by large errors, which can be expected to happen for ‘superstar’ players. Since these players are especially of interest to practitioners, the performance in the RMSE is considered the most important.
The losses are determined on different parts of the test set. First, the loss values of the RMSE and MAE are calculated using the complete test sets. Second, the loss function of the RMSE is also considered for different age groups on the test sets because estimating the potential of a player is mostly interesting for young players. Third, the RMSE is studied on important subgroups of players, like players with large positive or negative development, players with good performance, or players with a high transfer value. By determining the test losses separately on the general population of players, young players, and important subgroups of players, the models are studied on their predictive performance.
Beforehand, 5% of all football players are left out of both data sets based on stratified sampling for internal studies by the data provider. Because the data of player performance indicators is often dependent on time (Davis et al., 2024; Mendes-Neves et al., 2022), time-dependent train–test splits are applied to study the predictive performance on unseen data, as visualized in Figure 2. In both case studies, all data up to 2020 is considered as the training set, and from 2021 and later as the test set.
@@ 3.2.2 Model Training
In this study, linear, tree-based, and kNN-based models are implemented. The linear models are ordinary least squares (OLS), also known as multiple linear regression (MLR), lasso regression, and a linear mixed effect model (LME). For the OLS model (Seabold and Perktold, 2010), feature selection is performed by applying backwards selection with a threshold of the p-values of 0.0001 for the player quality and 0.001 for the player value. These values are smaller than commonly used for significance testing due to the predictive nature of this study and the large data sets combined with the fact that common assumptions of the linear regression model, like normality, do not hold. These p-values were selected via trial and error, considering values of the form $10^{k}$, such that roughly half of the features is excluded. For lasso regression (Pedregosa et al., 2011), feature selection is applied by selecting only the variables with nonzero coefficients. The linear mixed effect (LME) model (Seabold and Perktold, 2010) is trained to take into account the influence of a player’s nationality as a random influence. The feature selection for the LME model is performed by taking the 20 variables corresponding to the largest feature importances within the lasso model. This is conducted for computational feasibility. For these three linear models, no interaction effects are included.
@@
The tree-based models are the decision tree (Pedregosa et al., 2011), random forest (Pedregosa et al., 2011), and XGBoost (Chen and Guestrin, 2016) models. For the tree-based models, feature selection is performed by first adding noise variables to the data set and training the models. All features with a larger feature importance than the noise variables are then selected. As tree-based methods can have a bias favoring non-discrete variables, both discrete and continuous variables are added. The discrete features are compared with the discrete noise variables, and the continuous features with the continuous noise variables.
@@
To investigate the predictive power of time series, three k-nearest neighbors (kNN) models have been implemented using the Hierarchical Navigable Small Worlds indexer provided by Douze et al. (2024). This was carried out on an altered feature set consisting of time series information with lagged versions of the most important player indicators. Because the kNN model suffers from the curse of dimensionality (Section 2.5, Hastie et al. (2009)), the feature set is kept relatively small. The features and labels of this altered data set are described in Tables S3 and S6 in the Supplementary Materials. First, a normal kNN model is applied to the predictive problem. This model searches for the most similar data points with respect to the Euclidean norm. The prediction is then obtained by applying a weighted average, where closer data points are weighted more heavily as described by Dudani (1976). Possible weighting methods are the reciprocal of the absolute value of the distance, the distance with min–max scaling, or uniform weights. The method of calculating these weights is considered a hyperparameter. The second kNN model is constructed in the same way as the first, but it calculates the distances based on the Mahalanobis distance instead of the Euclidean distance. The Mahalanobis distance projects the features onto a decorrelated feature space and calculates the Euclidean distance in this changed feature space. This makes it possible to better distinguish differences because the lagged time series features are heavily dependent. Lastly, an adapted RReliefF method (Slifka and Whitton, 2003) is implemented to calculate feature importances of the normal kNN model in a regression context. The features are then multiplied by the feature importances before calculating Euclidean distances to introduce new feature weights. The kNN model with Euclidean distance is then trained on these reweighted features.
@@ 3.2.3 Hyperparameter Tuning
To select the best hyperparameters of the models in the case studies, Bayesian optimization (Head et al., 2021) is applied to find hyperparameters minimizing the RMSE loss. Because of the time dependencies of player performance indicators, cross-validation with an expanding windows split was implemented. This time series split strategy splits the data per year and incrementally grows the training set as visualized for the player quality prediction problem in Figure 3. After the hyperparameter tuning, the models are trained on the complete training set with the optimized hyperparameters.